### v1

# Bilinear spherical maximal function with fractal dilations

Abstract.
In this paper, we investigate $L^{p}$-boundedness of the bilinear spherical maximal function associated with a general dilation set $E\subset{\mathbb{R}}_{+}$. We quantify the range of $L^{p}$-boundedness in terms of a dilation-invariant notion of upper Minkowski dimension of the set $E$. A particular case of this study, settles an open question of $L^{p}$-boundedness of the lacunary bilinear spherical maximal function at borderline cases $p_{1}=1$ or $p_{2}=1$ in dimension $d\geq 4$.
2010 Mathematics Subject Classification:
Primary 42B15, 42B25
1. Introduction
Let $A_{t}f(x)$ denote the average of $f$ on the sphere of radius $t>0$ centered at $x\in{\mathbb{R}}^{d},~d\geq 2$. It is defined by
$A_{t}f(x)=\int_{\mathbb{S}^{d-1}}f(x-ty)d\sigma(y),$
where $d\sigma$ denotes the normalized surface measure on $\mathbb{S}^{d-1}.$ For a dilation set $E\subset{\mathbb{R}}_{+}$, the corresponding spherical maximal operator is defined by
$M_{E}f(x)=\sup_{t\in E}|A_{t}f(x)|.$
The $L^{p}$-bounds for the maximal operator $M_{E}$ have been extensively studied in the literature. Stein [SphStein] used Fourier analytic methods to obtain optimal $L^{p}$-bounds for the maximal function $M_{{\mathbb{R}}_{+}}f$ for $d\geq 3.$ The Fourier analytic approach does not extend to resolve the problem in dimension $d=2$. Later, Bourgain [CirBourgain] introduced some geometric agruments to settle the case of dimension $d=2$. We refer to the maximal function $M_{{\mathbb{R}}_{+}}f$ as the full spherical maximal function and denote it by $M_{full}f$. At the end-point $p=\frac{d}{d-1}$, Bourgain [sphBourgain] proved the restricted weak-type bounds in dimensions $d\geq 3$, while it was shown by Seeger, Tao and Wright [STW] that the restricted weak-type inequality fails in dimension $d=2$.
Under restrictions on the dilation set $E\subset{\mathbb{R}}_{+}$, the $L^{p}$-boundedness of $M_{E}$ extends to a larger range of $L^{p}$-spaces. The range of $p$ depends on the geometry and appropriate notion of size of the set $E$. For a general dilation set $E\subset{\mathbb{R}}_{+}$, Seeger, Wainger and Wright [SWW1] obtained sharp $L^{p}$-bounds for $p>p(E)$, where $p(E)$ depends on the dilation-invariant notion of the Minkowski dimension. The end-point cases were resolved in [STW]. A particularly interesting case arises when we consider $E$ to be a lacunary set. In this case, the $L^{p}$-bounds of the operator $M_{E}$ (denoted by $M_{lac}$), holds for all $p>1$, see [lacsphCalderon, CoifmanWeiss, RubioDuoandi]. It is conjectured that the lacunary maximal function $M_{lac}$ is of weak-type $(1,1)$, however the problem remains open till date. We refer to  [CladekKrause] for recent developments in this direction. The sharp range of sparse bounds for $M_{full}$ and $M_{lac}$ was obtained by Lacey [sparseLacey] using continuity estimates and $L^{p}$-improving bounds.
When $E\subset[1,2]$, the corresponding maximal operator $M_{E}$ is referred to as the local maximal operator. Due to the local nature of the operator, it satisfies improving $L^{p}\to L^{q}$-bounds for some $q\geq p$. Such estimates for the local operator $M_{E}$, in terms of the upper Minkowski dimension and the Assouad spectrum of $E$, have been studied in [AHRS, RoosSeeger, LRZZ]. In these results, the Fourier decay of surface measures plays a key role, and the fractal geometry of $E$ is captured through the covering numbers and Minkowski dimension.
Let us recall the definition of dilation-invariant notion of the Minkowski dimension, see [SWW1] for more details.
Definition 1.1.
Let $E\subset{\mathbb{R}}_{+}$ and $E_{k}=E\cap[2^{-k},2^{-k+1}]$, $k\in{\mathbb{Z}}$. We say $E$ have upper Minkowski dimension $\beta$ if
$\sup_{\delta>0}\sup_{k\in{\mathbb{Z}}}\frac{\log N(E_{k},2^{-k}\delta)}{\log\delta^{-1}}=\beta,$
where $N(T,\delta)$ denotes the minimum number of $\delta$-length intervals needed to cover $T$.
Note that the definition of upper Minkowski dimension is uniform in dyadic scales.
1.1. Bilinear spherical maximal function
The bilinear spherical maximal function associated to a given dilation set $E\subset{\mathbb{R}}_{+}$ is defined by
$\mathcal{M}_{E}(f_{1},f_{2})(x):=\sup_{t\in E}|\mathcal{A}_{t}(f_{1},f_{2})(x)|,$
where the bilinear spherical average $\mathcal{A}_{t}(f_{1},f_{2})(x)$ is defined by
$\mathcal{A}_{t}(f_{1},f_{2})(x)=\int_{\mathbb{S}^{2d-1}}f_{1}(x-ty_{1})f_{2}(x-ty_{2})\;d\sigma(y_{1},y_{2}).$
The bilinear spherical averages first appeared in the work [GGIPS]. The $L^{p}$-estimates for the full maximal operator $\mathcal{M}_{full}:=\mathcal{M}_{{\mathbb{R}}_{+}}$ were studied in [BGHHO, GHH, HHY]. The sharp range of $L^{p}$-boundedness, including the end-point cases, were obtained by Jeong and Lee [JeongLee] for dimensions $d\geq 2$. They introduced the method of slicing and gave a simple proof of optimal $L^{p}$-estimates for the operator $\mathcal{M}_{full}$. The slicing argument provides with a decomposition of $\mathbb{S}^{2d-1}$ into a family of lower dimensional spheres. More precisely, Jeong and Lee [JeongLee] showed that if $F$ is a continuous function defined on ${\mathbb{R}}^{2d},d\geq 2$ and $(y,z)\in{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$, then
$\int_{\mathbb{S}^{2d-1}}F(y,z)\;d\sigma(y,z)=\int_{B^{d}(0,1)}\int_{\mathbb{S}^{d-1}}F(y,\sqrt{1-|y|^{2}}z)\;d\sigma_{d-1}(z)(1-|y|^{2})^{\frac{d-2}{2}}dy.$ (1.1)
The formula as above yields the following pointwise estimate.
$\mathcal{M}_{full}(f,g)(x)\lesssim M_{HL}f(x)M_{full}g(x),$ (1.2)
where $M_{HL}f$ is the Hardy-Littlewood maximal function. Using the symmetry of the bilinear spherical maximal function in $f$ and $g$, the boundedness region for the operator $\mathcal{M}_{full}$ is determined by using $L^{p}$-estimates of Hardy-Littlewood maximal function $M_{HL}f(x)$ and spherical maximal function $M_{full}g(x)$ for $d\geq 2.$ In dimension $d=1$, the optimal range of boundedness of $\mathcal{M}_{full}$ was obtained in [ChristZhou, DosidisRamos, BCSS].
Jeong and Lee [JeongLee] also obtained $L^{p}$-improving estimates for the local version of bilinear spherical maximal function using the slicing method. Later, in [BCSS], the authors improved the range of $L^{p}$-improving estimates for the local bilinear spherical maximal function. Observe that for a general dilation set $E$, the slicing argument always leads to pointwise domination of the bilinear maximal function $\mathcal{M}_{E}(f,g)$ by a product of the Hardy-Littlewood maximal function and the full spherical maximal function as in the estimate (1.2). This does not improve the range of $L^{p}$-boundedness of the operator $\mathcal{M}_{E}$. Therefore, the slicing argument has its limitations in dealing with the bilinear maximal function $\mathcal{M}_{E}(f,g)$ associated with a dilation set $E$.
The $L^{p}$-estimates for the bilinear lacunary spherical maximal operator $\mathcal{M}_{lac}=:\mathcal{M}_{E}$, where $E$ is a lacunary set, were obtained by Borges and Foster [BorgesFoster] for dimensions $d\geq 2$, and by Christ and Zhou  [ChristZhou] for dimension $d=1$. The proof for dimensions $d\geq 2$ relies mainly on an $L^{2}\times L^{2}\rightarrow L^{1}$-decay estimate in the frequency parameter for single scale average $\mathcal{A}_{1}(f,g)$, where functions $f$ or $g$ are frequency localized in an appropriate annular region formed by Littlewood-Paley decomposition. Then, the desired estimates for the bilinear lacunary maximal function are obtained by a boot-strapping type argument. The decay estimate at the point $(2,2,1)$ is obtained by a combination of slicing argument (1.1) and Fourier decay property of $\widehat{d\sigma}$. Due to unavailability of the slicing method and lack of appropriate Fourier decay of $\widehat{d\sigma}$ in dimension $d=1$, it is significantly more difficult to obtain $L^{p}$-estimates for the bilinear lacunary maximal function in dimension $d=1$. Christ and Zhou  [ChristZhou] overcame these issues and proved trilinear smoothing estimates for bilinear averaging operators along certain curves, including the averages along the circle. The trilinear smoothing estimates play a key role in obtaining decay estimates for frequency localized pieces and finally obtaining the desired $L^{p}$-estimates for $\mathcal{M}_{lac}$. Here we summarize the results for the bilinear lacunary maximal operator $\mathcal{M}_{lac}$.
Theorem A.
[ChristZhou, BorgesFoster] Let $1\leq p_{1},p_{2}\leq\infty$ with $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}$. Then, $\mathcal{M}_{lac}$ maps $L^{p_{1}}(\mathbb{R}^{d})\times L^{p_{2}}(\mathbb{R}^{d})$ to $L^{p}(\mathbb{R}^{d})$ if
  1. (i)
    $d=1$ and $1<p_{1},p_{2}\leq\infty$.
  2. (ii)
    $d\geq 2$ and $1<p_{1},p_{2}\leq\infty$, or $p_{1}=1$ and $p_{2}\in(\frac{d}{d-1},\infty)$, or $p_{2}=1$ and $p_{1}\in(\frac{d}{d-1},\infty)$.
Moreover, the weak-type bounds $\mathcal{M}_{lac}:L^{1}(\mathbb{R}^{d})\times L^{\infty}(\mathbb{R}^{d})\rightarrow L^{1,\infty}(\mathbb{R}^{d})$ and $\mathcal{M}_{lac}:L^{\infty}(\mathbb{R}^{d})\times L^{1}(\mathbb{R}^{d})\rightarrow L^{1,\infty}(\mathbb{R}^{d})$ hold.
[…]
Figure 1. Boundedness regions for the operator $\mathcal{M}_{lac}$ in A.
Christ and Zhou [ChristZhou] showed that the strong-type $L^{p}$-bounds for $\mathcal{M}_{lac}$ do not hold if either $p_{1}=1$ or $p_{2}=1$. However, in higher dimensions, $L^{p}$-boundedness of $\mathcal{M}_{lac}$ is known to hold for the range $p_{1}=1$ and $p_{2}\in(\frac{d}{d-1},\infty)$, or $p_{2}=1$ and $p_{1}\in(\frac{d}{d-1},\infty)$. The question concerning the remaining points on the line segments $p_{1}=1$ and $p_{2}=1$ remained unresolved. In this paper, we address this problem and extend the range of $L^{p}$-boundedness of $\mathcal{M}_{lac}$ to the borderline cases in $d\geq 4$ excepting the end-point point $(1,1)$. Indeed, we obtain new results for bilinear spherical maximal function associated with general dilation sets. The study of local bilinear spherical maximal function with a general dilation set $E\subset[1,2]$ was carried out by Borges, Foster and Ou [BFO]. They proved Sobolev smoothing estimates at the point $(2,2,2)$ for the local maximal operator along hypersurfaces that exhibit Fourier decay of order $\frac{1}{(1+|\xi|)^{\alpha}}$ with $\alpha>\frac{d}{2}$. They also obtained necessary conditions for the boundedness of $\mathcal{M}_{E}$ in terms of the upper Minkowski dimension of the underlying fractal dilation set. The Sobolev smoothing estimates give a simplified derivation of the continuity estimates for such local maximal operators associated with general dilation sets $E$. They also deduced $L^{p}$-improving bounds and sparse bounds for the associated bilinear maximal operators. However, when restricted to spherical maximal functions, this does not improve the $L^{p}$-boundedness range for $\mathcal{M}_{E}$ beyond that of $\mathcal{M}_{full}$.
The sparse bounds for the bilinear spherical maximal function $\mathcal{M}_{full}$ were studied by Borges et al. in [BFOPZ]. Extensions of these results to some general surfaces have also been explored; for instance, see [CGHHS, LS] for $L^{p}$-estimates and [RSS, PS] for sparse bounds. Additionally, the multilinear extension of the maximal function associated with spheres and other general surfaces has been studied in [Dosidis, SS, GHHP, CLS, Gao].
2. Main results
In this paper, we explore alternate of slicing method to obtain new results concerning $L^{p}$-boundedness of the bilinear maximal operator $\mathcal{M}_{E}$ associated with general dilation set $E$. We intend to quantify the range of exponents $(p_{1},p_{2},p)$ for the $L^{p}$-boundedness of the operator $\mathcal{M}_{E}$ in terms of upper Minkowski dimension of the set $E$. We exploit the Fourier decay property of the surface measure in a refined manner and perform a scale-adapted discretization of $E$ to capture the fractal geometry of the set $E$. These ideas, along with suitable adaptation of bilinear Calderón-Zygmund theory, yield new results for the maximal operator $\mathcal{M}_{E}$. In order to describe our results, we need the following notation.
Let $\Omega({P_{1},P_{2},\dots,P_{k}})$ denote the convex hull of points $\{P_{1},P_{2},\dots,P_{k}\}\subset{\mathbb{R}}^{2}.$ Let $d\geq 2$ and set $O=(0,0),\;A=(1,0),\;B=(0,1),\;P_{0}=\big(\frac{2d-1}{2d},\frac{2d-1}{2d}\big),\;Q_{0}=\big(1,\frac{d-1}{d}\big),\;\text{and}~R_{0}=\big(\frac{d-1}{d},1\big)$. For $0\leq\beta\leq 1$ consider the point
$P=P(d,\beta)=\begin{cases}\left(\frac{2d-2}{2d-2+\beta},\frac{2d-2}{2d-2+\beta}\right)&\beta<\frac{2d-2}{2d-1},\\ P_{0}&\beta\geq\frac{2d-2}{2d-1}.\end{cases}$
Further, for $d\geq 4$, we define
$Q=Q(d,\beta)=\begin{cases}\left(1,\frac{d-3}{d-3+\beta}\right)&\beta<\frac{d-3}{d-1},\\ Q_{0}&\beta\geq\frac{d-3}{d-1},\end{cases}\qquad\text{and}\qquad R=R(d,\beta)=\begin{cases}\left(\frac{d-3}{d-3+\beta},1\right)&\beta<\frac{d-3}{d-1},\\ R_{0}&\beta\geq\frac{d-3}{d-1}.\end{cases}$
Throughout this paper, we shall assume that exponents $(p_{1},p_{2},p)$ satisfying $1\leq p_{1},p_{2}\leq\infty$ and the Hölder condition $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}$ unless specified otherwise. We also use the notation $A\lesssim B$ to denote that $A\leq CB$ for some $C>0$.
The following is the main result of the paper.
Theorem 2.1.
Let $d\geq 2$ and $0<p_{1},p_{2},p\leq\infty$ with $\frac{1}{p_{1}}+\frac{1}{p_{2}}=\frac{1}{p}$. Assume that $E\subset{\mathbb{R}}_{+}$ has an upper Minkowski dimension $\beta$. Then the operator $\mathcal{M}_{E}$ is bounded from $L^{p_{1}}(\mathbb{R}^{d})\times L^{p_{2}}(\mathbb{R}^{d})$ into $L^{p}(\mathbb{R}^{d}),$ where
  1. (1)
    $d=2,3$, and $(\frac{1}{p_{1}},\frac{1}{p_{2}})\in\Omega(\{O,A,Q_{0},P,R_{0},B\})\setminus\{A,B,PQ_{0},PR_{0}\}$.
  2. (2)
    $d\geq 4$, and $(\frac{1}{p_{1}},\frac{1}{p_{2}})\in\Omega(\{O,A,Q,P,R,B\})\setminus\{A,B,PQ,PR\}$.
Moreover, if the set $E\subseteq[1,2]$ has upper Minkowski dimension $\beta$, then the local bilinear maximal function $\mathcal{M}_{E}$ satisfies restricted weak-type estimate $L^{p_{1},1}(\mathbb{R}^{d})\times L^{p_{2},1}(\mathbb{R}^{d})\rightarrow L^{p,\infty}(\mathbb{R}^{d})$ whenever $\Big(\frac{1}{p_{1}},\frac{1}{p_{2}}\Big)$ lies on the line segments $QP$ or $PR$.
[…]
Figure 2. The dark gray region depicts the improved range for boundedness of $\mathcal{M}_{E}$.
The lacunary bilinear maximal function corresponds to the case $\beta=0$ in the theorem above. Therefore, we have $L^{p}$-boundedness of the lacunary bilinear spherical maximal operator $\mathcal{M}_{lac}$ in the borderline cases in dimension $d\geq 4.$
Theorem 2.2.
[…]
Figure 3. Boundedness region for the operator $\mathcal{M}_{lac}$ in dimension $d\geq 4$.
Moreover, due to Stein’s maximal principle, we have the following result about pointwise convergence of bilinear spherical averages over an arbitrary set of dilations.
Corollary 2.3.
Let $\{t_{j}\}$ be a sequence of positive real numbers such that $t_{j}\rightarrow 0$. Let $E=\{t_{j}:j\in{\mathbb{N}}\}$ have the upper Minkowski dimension $\beta$. Then for $f_{1}\in L^{p_{1}}({\mathbb{R}}^{d})$ and $f_{2}\in L^{p_{2}}({\mathbb{R}}^{d})$, where $\big(\frac{1}{p_{1}},\frac{1}{p_{2}}\big)$ lies in the region given in Theorem 2.1, the following pointwise almost everywhere convergence holds
$\lim_{j\to\infty}\mathcal{A}_{t_{j}}(f_{1},f_{2})(x)=f_{1}(x)f_{2}(x).$
2.1. Strategy of proof
As indicated earlier, the slicing method does not improve the range of $L^{p}$-estimates for the bilinear spherical maximal functions associated with general dilation sets beyond that of the full maximal operator $\mathcal{M}_{full}$. We follow the following strategy to prove Theorem 2.1.
  • First, we perform Littlewood-Paley decomposition of both the functions in frequency space. The $L^{p}$-estimates for the maximal function acting on low frequency terms and terms with mixed type frequencies can be dealt with directly. In this case, we could dominate the maximal function by a product of Hardy-Littlewood maximal functions, see Lemma 3.1. This step reduces the problem to proving the desired bounds for terms with high frequencies.
  • Next, we consider the local maximal operator associated with the set $E_{0}=E\cap[1,2]$. In Section 4 we carry out the decomposition of multiplier as used by Heo, Hong and Yang  [HHY].
  • We derive
    $L^{2}\times L^{2}\to L^{1},\quad L^{2}\times L^{1}\to L^{2/3},\quad\text{and}~L^{1}\times L^{2}\to L^{2/3}$
    estimates, see Lemma 5.2 with suitable decay for the operator norm depending on the frequency parameter and the upper Minkowski content of $E_{0}$. The second estimate that we obtain is an end-point estimate for the operator. We use an appropriate space localization of the operator and establish end-point $L^{1}\times L^{1}\to L^{1/2}$-estimate, see Lemma 5.3. This end-point estimate has growth in terms of the frequency parameter and the number of intervals used in discretization of the $E_{0}$.
  • The interpolation argument between the growth estimate at end-point and the decay estimate yields a region of $L^{p}$-boundedness of $\mathcal{M}_{E_{0}}$ depending on upper Minkowski dimension of $E_{0}$.
  • Finally, in Section 6, we extend the results obtained for local maximal operator to the global operators associated with $E\subset{\mathbb{R}}_{+}$. This part is done by using appropriate rescaling arguments and adaptation of the bilinear Calderón-Zygmund theory to the underlying situation. We exploit the methods used in [ChristZhou] and [BCSS2].
3. Reduction using Littlewood-Paley decomposition
Consider the following partition of the identity
$\widehat{\phi}(\xi)+\sum_{j=1}^{\infty}\widehat{\psi}_{2^{-j}}(\xi)=1,\quad\xi\neq 0,$
where $\phi\in\mathcal{S}({\mathbb{R}}^{d})$ is a radial and decreasing function such that $\text{supp}(\widehat{\phi})\subseteq B(0,2)$ and $\widehat{\phi}(\xi)=1$ for $\xi\in B(0,1)$. We have used the notation $\widehat{\phi_{t}}(\xi)=\widehat{\phi}(t\xi)$ and $\widehat{\psi}_{t}(\xi)=\widehat{\phi}(t\xi)-\widehat{\phi}(2t\xi)$.
For $i\in{\mathbb{Z}}$, we define smooth Fourier projection operators $\widehat{P_{i}f}(\xi)=\hat{f}(\xi)\hat{\phi}(2^{-i}\xi)$ and $\widehat{R_{i}f}=\hat{f}(\xi)\hat{\psi}(2^{-i}\xi)$. Using this, we can write
$f=P_{i}f+\sum_{j=1}^{\infty}R_{i+j}f.$
Consequently, we get the following decomposition of the operator $\mathcal{M}_{E}$.
$\mathcal{M}_{E}(f_{1},f_{2})(x)$
$\leq$ $\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(P_{k}f_{1},P_{k}f_{2})(x)\Big|+\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\bigg|\mathcal{A}_{t}\Big(\sum^{\infty}_{n_{1}=1}R_{k+n_{1}}f_{1},P_{k}f_{2}\Big)(x)\bigg|$
$+\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\bigg|\mathcal{A}_{t}\Big(P_{k}f_{1},\sum^{\infty}_{n_{2}=1}R_{k+n_{2}}f_{2}\Big)(x)\bigg|+\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}\Big(\sum^{\infty}_{n_{1}=1}R_{k+n_{1}}f_{1},\sum^{\infty}_{n_{2}=1}R_{k+n_{2}}f_{2})(x)\Big)\bigg|$
$\lesssim$ $\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(P_{k}f_{1},P_{k}f_{2})(x)\Big|+\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(P_{k}f_{1},f_{2})(x)\Big|$
$+\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(f_{1},P_{k}f_{2})(x)\Big|+\sum^{\infty}_{n_{1},n_{2}=1}\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x),$ (3.1)
where
$\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x):=\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(R_{k+n_{1}}f_{1},R_{k+n_{2}}f_{2})(x)\Big|.$
Denote
$\mathcal{M}^{\mathbf{n}}_{k}(f_{1},f_{2})(x)=\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(R_{k+n_{1}}f_{1},R_{k+n_{2}}f_{2})(x)\Big|.$
The first three terms in (3) are controlled by product of the Hardy-Littlewood maximal functions, which in turn implies strong-type $L^{p}$-boundedness of $\mathcal{M}_{E}$ for $p_{1},p_{2}>1$ and weak-type $L^{p}$-boundedness when $p_{i}=1$. The strong-type $L^{p}$-boundedness of the operator $\mathcal{M}_{E}$ in the case where $p_{i}=1$ and $1<p_{j}<\infty$ for $i,j\in\{1,2\},~i\neq j$, are obtained by standard interpolation argument between weak-type bounds at $\big(1,1,\frac{1}{2}\big)$ and either at $(1,\infty,1)$ or $(\infty,1,1)$.
Lemma 3.1.
Let $d\geq 2$. The following pointwise estimates hold.
$\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(P_{k}f_{1},f_{2})(x)\Big|$ $\lesssim M_{HL}f_{1}(x)M_{HL}f_{2}(x),$
$\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(f_{1},P_{k}f_{2})(x)\Big|$ $\lesssim M_{HL}f_{1}(x)M_{HL}f_{2}(x),$
$\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(P_{k}f_{1},P_{k}f_{2})(x)\Big|$ $\lesssim M_{HL}f_{1}(x)M_{HL}f_{2}(x).$
Proof.
Note that if $t\in E_{k}$ and $|y|\leq 1$, then we have $|ty|\leq 2^{-k+1}$. Since $\phi\in\mathcal{S}({\mathbb{R}}^{d})$, we get
$\sup_{t\in E_{k}}\sup_{|y|\leq 1}|P_{k}f_{1}(x-ty)|$ $\leq\sup_{|ty|\leq 2^{-k+1}}\left|\int_{{\mathbb{R}}^{d}}f_{1}(z)2^{kd}\phi(2^{k}(x-ty-z))~dz\right|$
$\lesssim\sup_{|y|\leq 2}\left|\int_{{\mathbb{R}}^{d}}\frac{2^{kd}f_{1}(z)}{\Big(4+|2^{k}(x-z)-y|\Big)^{N}}dz\right|$
$\lesssim M_{HL}f_{1}(x).$
Therefore, using the slicing argument, we have
$\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(P_{k}f_{1},f_{2})(x)\Big|$ $\leq\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\left(\sup_{|y|\leq 1}|P_{k}f_{1}(x-ty)|\int_{\mathbb{S}^{2d-1}}|f_{2}(x-tz)|d\sigma(y,z)\right)$
$\lesssim M_{HL}f_{1}(x)\sup_{k\in{\mathbb{Z}}}\sup_{t\in E_{k}}\int_{B^{d}(0,1)}|f_{2}(x-tz)|(1-|z|^{2})^{(d-2)/2}\int_{\mathbb{S}^{d-1}}d\sigma(y)~dz$
$\lesssim M_{HL}f_{1}(x)M_{HL}f_{2}(x).$
The remaining two estimates are deduced similarly. ∎
In view of the lemma above, we are left with proving the desired $L^{p}$-boundedness of the operator $\mathcal{M}^{\mathbf{n}}$, with a suitbale decay in $|\mathbf{n}|$, in order to complete the proof of Theorem 2.1.
4. Decomposition of multiplier
Consider the multiplier form of the bilinear spherical average given by
$\mathcal{A}_{t}(f_{1},f_{2})(x)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{f_{1}}(\xi)\widehat{f_{2}}(\eta)\widehat{d\sigma}(t\xi,t\eta)e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,$
where $\widehat{d\sigma}$ denotes the Fourier transform of the spherical measure and has the expression
$\widehat{d\sigma}(t\xi,t\eta)=\frac{2\pi J_{d-1}(|(t\xi,t\eta)|)}{|(t\xi,t\eta)|^{d-1}}.$ (4.1)
Here, $J_{d-1}$ denotes the Bessel function of order d1d-1. The following asymptotic estimate of the Bessel function holds
$J_{d-1}(r)\sim r^{-\frac{1}{2}}\bigg(e^{2\pi ir}\sum_{j=0}^{\infty}a_{j}r^{-j}+e^{-2\pi ir}\sum_{j=0}^{\infty}a_{j}r^{-j}\bigg)\quad\text{ as }r\to\infty,$
for some coefficients $a_{j}$ and $b_{j}$, such that for all nonnegative integers $N$ and $\alpha$
$\bigg(\frac{d}{dr}\bigg)^{\alpha}\Bigg[J_{d-1}(r)-r^{-\frac{1}{2}}\bigg(e^{2\pi ir}\sum_{j=0}^{N}a_{j}r^{-j}+e^{-2\pi ir}\sum_{j=0}^{N}a_{j}r^{-j}\bigg)\Bigg]=O\big(r^{-\alpha-\frac{N+1}{2}}\big),$ (4.2)
as $r\to\infty$. We refer to [[book-Stein], Chapter VIII] for details.
In view of the asymptotic estimates of Bessel functions, observe that the main term in the multiplier corresponding to the operator $\mathcal{A}_{t}$ behaves like $\frac{e^{2\pi i|(t\xi,t\eta)|}}{|(t\xi,t\eta)|^{\frac{2d-1}{2}}}.$ The remaining terms satisfy higher order decay and hence we need to focus on the main term only. Therefore, we prove the desired $L^{p}$-boundedness of the maximal operator corresponding to bilinear operators defined by
$T^{\mathbf{n}}_{t,k}(f_{1},f_{2})(x)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{R_{k+n_{1}}f_{1}}(\xi)\widehat{R_{k+n_{2}}f_{2}}(\eta)m(t\xi,t\eta)e^{2\pi it|(\xi,\eta)|}e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,$
where $t\in E_{k}$ and $m(\xi,\eta)=|(\xi,\eta)|^{-\frac{2d-1}{2}}$.
As described previously, the $L^{p}$-estimates for the global maximal operator are obtained via suitable adaptation of bilinear Calderón-Zygmund theory. We shall restrict ourselves to the case of unit scale, that is $k=0$. The proof is easily adapted to obtain the corresponding estimates for general dyadic scale $2^{k}$.
Let us use the notation $T^{\mathbf{n}}_{t}=T^{\mathbf{n}}_{t,0},$ where $t\in E_{0}$. We fix $t\in E_{0}$ and work with the operator $T^{\mathbf{n}}_{t}.$ Due to the lack of additional decay in $|(\xi,\eta)|$ for the derivative of the term $e^{2\pi it|(\xi,\eta)|},$ we need to decompose this exponential term further.
Let $N$ be a positive integer such that $N\sim 2^{n_{1}}.$ Write the exponential term as follows
$e^{2\pi it|(\xi,\eta)|}=e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi it\sqrt{N^{2}+|\eta|^{2}}},$
where
$\Phi_{N}(\xi,\eta):=|(\xi,\eta)|-\sqrt{N^{2}+|\eta|^{2}}=\frac{(|\xi|-N)(|\xi|+N)}{\sqrt{|\xi|^{2}+|\eta|^{2}}+\sqrt{N^{2}+|\eta|^{2}}}.$
Note that $\Phi_{N}$ satisfies the following estimate
$|\partial^{\alpha}_{\eta}\Phi_{N}(\xi,\eta)|\lesssim|\eta|^{-|\alpha|},$ (4.3)
for all multi-indices $\alpha$ uniformly in $\xi$ when $N\leq|\xi|<N+1$.
Let $\chi_{{}_{N}}$ denote the characteristic function of the interval $[N,N+1)$. Define
$F_{N}f(x):=\int_{\mathbb{R}^{d}}\widehat{f}(\xi)\chi_{{}_{N}}(|\xi|)e^{2\pi ix\cdot\xi}d\xi.$
This gives us the following decomposition of the operator $T^{\mathbf{n}}_{t}.$
$T^{\mathbf{n}}_{t}(f_{1},f_{2})(x)$ $=\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x),$
where
$T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}m_{t}^{\mathbf{n}}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\widehat{F_{N}f_{1}}(\xi)\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta)e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,$
$\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta):=\widehat{f_{2}}(\eta)\widehat{\psi}\left(2^{-n_{2}}\eta\right)e^{2\pi it\left(\sqrt{N^{2}+|\eta|^{2}}\right)},$ (4.4)
and
$m^{\mathbf{n}}_{t}(\xi,\eta):=m(t\xi,t\eta)\widehat{\psi}\left(2^{-n_{1}}\xi\right)\widehat{\zeta}\left(2^{-n_{2}}\eta\right).$
Here $\zeta\in\mathcal{S}({\mathbb{R}}^{d})$ such that $\widehat{\zeta}$ is supported in a small neighbourhood of $\operatorname{supp}(\psi)$ and $\widehat{\zeta}(\eta)=1$ for $\eta\in\operatorname{supp}(\psi)$.
Using the estimate (4.3), we get that
$|\partial^{\alpha}_{\eta}\big(m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\big)|\lesssim C_{\alpha}2^{-\frac{2d-1}{2}|\mathbf{n}|}2^{-|\alpha|n_{2}},$ (4.5)
for all multi-indices $\alpha$ uniformly in $t\in E_{0}$.
Since $\widehat{U_{t,N}^{n_{2}}f_{2}}(\eta)=\int_{{\mathbb{R}}^{d}}U_{t,N}^{n_{2}}f_{2}(y)e^{-2\pi i\eta\cdot y}dy$, we can rewrite
$T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{F_{N}f_{1}}(\xi)U_{t,N}^{n_{2}}f_{2}(y)\Big(\int_{{\mathbb{R}}^{d}}m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi i\eta\cdot(x-y)}d\eta\Big)e^{2\pi ix\cdot\xi}\;d\xi dy.$ (4.6)
Using integration by parts argument, we obtain
$\int_{{\mathbb{R}}^{d}}m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi i\eta\cdot(x-y)}d\eta$ (4.7)
$=\frac{1}{(1+4\pi^{2}|2^{n_{2}}(x-y)|^{2})^{\lceil\frac{d}{2}\rceil+1}}\int_{{\mathbb{R}}^{d}}(1-2^{2n_{2}}\Delta_{\eta})^{\lceil\frac{d}{2}\rceil+1}\Big(m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\Big)e^{2\pi i\eta\cdot(x-y)}d\eta,$
where we have used the fact that
$(1-2^{2n_{2}}\Delta_{\eta})e^{2\pi i\eta\cdot(x-y)}=(1+4\pi^{2}|2^{n_{2}}(x-y)|^{2})e^{2\pi i\eta\cdot(x-y)}.$
Substituting (4.7) into (4.6), we get
$T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)$ $=\int_{\mathbb{R}^{d}}A_{n_{2}}\left(U_{t,N}^{n_{2}}f_{2}\right)(x,\eta)B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta)d\eta,$
where
$B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta):=$ $\int_{\mathbb{R}^{d}}\left(1-2^{2n_{2}}\Delta_{\eta}\right)^{\lceil\frac{d}{2}\rceil+1}\left(m^{\mathbf{n}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\right)\widehat{F_{N}f_{1}}(\xi)e^{2\pi ix\cdot\xi}d\xi,$ (4.8)
and
$A_{n_{2}}\left(U^{n_{2}}_{t,N}f_{2}\right)(x,\eta):=\int_{\mathbb{R}^{d}}\frac{U^{n_{2}}_{t,N}f_{2}(y)e^{2\pi i\eta\cdot(x-y)}}{\left(1+4\pi^{2}\left|2^{n_{2}}(x-y)\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy.$
Note that
$|A_{n_{2}}\left(U^{n_{2}}_{t,N}f_{2}\right)(x,\eta)|$ $\leq\int_{\mathbb{R}^{d}}\frac{|U^{n_{2}}_{t,N}f_{2}(y)|}{\left(1+4\pi^{2}\left|2^{n_{2}}(x-y)\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
$:=L_{n_{2}}\left(U_{t,N}^{n_{2}}f_{2}\right)(x).$
This gives us the following bound.
$|T^{\mathbf{n}}_{t,N}(f_{1},f_{2})(x)|\leq L_{n_{2}}\left(U_{t,N}^{n_{2}}f_{2}\right)(x)\int_{\mathbb{R}^{d}}\left|B^{\mathbf{n}}_{t,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta.$ (4.9)
We have the following $L^{p}$-estimates for the terms associated with operators $U_{t,N}^{n_{2}}$ and $B^{\mathbf{n}}_{t,N}$ introduced as above.
Lemma 4.1.
Let $a>0$ and $m\in C^{d+2}({\mathbb{R}}^{2d}\setminus\{(0,\dots,0)\}$ be such that
$|\partial_{\xi}^{\alpha}\partial_{\eta}^{\beta}m(\xi,\eta)|\lesssim_{\alpha,\beta}|(\xi,\eta)|^{-a}|\xi|^{-\alpha}|\eta|^{-\beta},$
for all multi-indices $|\alpha|+|\beta|\leq d+2$. For $\mathbf{n}\in{\mathbb{N}}^{2}$ and $N>0$, let $B^{\mathbf{n}}_{s,N}(f)$ be defined as in 4.8. Then, we have
$\left\|\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(f\right)(\cdot,\eta)\right|d\eta\right\|_{L^{2}}\lesssim 2^{-|\mathbf{n}|a}2^{n_{2}d}\|f\|_{L^{2}}.$
Lemma 4.2 ([HHY], Lemma 2.2).
Let $I\subset[1,2]$. For $n_{2}\in{\mathbb{N}}$ and $N>0$, let $U_{s,N}^{n_{2}}g$ be defined as in 4.4. Then, we have
$\left\|\bigg(\int_{I}\big|U_{s,N}^{n_{2}}g\big|^{2}ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\lesssim|I|^{\frac{1}{2}}2^{n_{2}\frac{d}{2}}\|g\|_{L^{1}}.$
The proof of Lemma 4.1 follows by imitating the proof of estimate (2.35) from [HHY] on pages 418-419. Lemma 4.2 was proved for $I=[1,2]$ in [HHY]. The same argument can be adapted for any interval $I\subset[1,2]$ to obtain Lemma 4.2. We skip the details here.
5. Estimate for local maximal operator
We use Bourgain’s interpolation argument to prove restricted weak-type estimates for the local maximal function $\mathcal{M}_{E_{0}}.$ We state the interpolation lemma for convenience. The reader is referred to [Lee1] for more details.
Lemma 5.1 ([Lee1], Lemma 2.6).
Let $\epsilon_{1},\epsilon_{2}>0$. Suppose that $\{T_{j}\}$ is a sequence of linear (or sublinear) operators such that for some $1\leq p_{1},p_{2}<\infty$, and $1\leq q_{1},q_{2}<\infty$,
$\|T_{j}(f)\|_{L^{q_{1}}}\leq M_{1}2^{\epsilon_{1}j}\|f\|_{L^{p_{1}}},~~\|T_{j}(f)\|_{L^{q_{2}}}\leq M_{2}2^{-\epsilon_{2}j}\|f\|_{L^{p_{2}}}.$
Then $T=\sum_{j}T_{j}$ is bounded from $L^{p,1}$ to $L^{q,\infty}$ with
$\|T(f)\|_{L^{q,\infty}}\lesssim M^{\theta}_{1}M^{1-\theta}_{2}\|f\|_{L^{p,1}},$
where $\theta=\frac{\epsilon_{2}}{\epsilon_{1}+\epsilon_{2}}$, $\frac{1}{q}=\frac{\theta}{q_{1}}+\frac{1-\theta}{q_{2}}$ and $\frac{1}{p}=\frac{\theta}{p_{1}}+\frac{1-\theta}{p_{2}}$.
Note that in view of the reduction step using Littlewood-Paley decomposition, see Section 3 and Lemma 3.1, we need to prove desired estimates for the maximal operator
$\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})(x)=\sup_{t\in E_{0}}\Big|\mathcal{A}_{t}(R_{n_{1}}f_{1},R_{n_{2}}f_{2})(x)\Big|.$
We have the following results for $\mathcal{M}^{\mathbf{n}}_{0}$.
Lemma 5.2.
Let $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
  1. (1)
    $\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}},$
  2. (2)
    $\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}},$
  3. (3)
    $\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
Lemma 5.3.
Let $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
$\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{1/2}}\lesssim 2^{|\mathbf{n}|\beta}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{1}}.$
Note that if we assume that Lemma 5.2 and Lemma 5.3 hold true, then bilinear analogue of real interpolation theorem, see Theorem 7.2.2 in [modernGrafakos], yields strong-type $L^{p}$-estimates
$\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{p}}\lesssim 2^{-|\mathbf{n}|\epsilon}\|f_{1}\|_{L^{p_{1}}}\|f_{2}\|_{L^{p_{2}}},$
for some $\epsilon>0$ and $\big(\frac{1}{p_{1}},\frac{1}{p_{2}}\big)\in\Omega(\{O,A,Q,P,R,B\})\setminus\{A,B,PQ,PR\}$. Therefore, summing over $n_{1}$ and $n_{2},$ we get the desired strong-type $L^{p}$-estimates for the operator $\mathcal{M}_{E_{0}}$.
Moreover, the bilinear complex interpolation theorem applied to estimates $(1)$ and $(2)$ of Lemma 5.2 gives us
$\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{\frac{2p}{p+2}}}\lesssim 2^{-|\mathbf{n}|\big(\frac{3d-1-\beta}{2}-\frac{d+1}{p}\big)}\|f_{1}\|_{L^{p}}\|f_{2}\|_{L^{2}},\quad 1\leq p\leq 2.$
Finally, for the operator $\mathcal{M}_{E_{0}}$, the restricted weak-type inequality at the point $\big(\frac{1}{p_{1}},\frac{1}{p_{2}}\big)$ lying on the line segment $PQ$ follows by Bourgain’s interpolation Lemma 5.1, applied to the preceding estimate and Lemma 5.3. The same argument could be used to obtain restricted weak-type estimates for points on the line segment $PR$. ∎
Therefore, we are left with proving Lemma 5.2 and Lemma 5.3.
5.1. Proof of Lemma 5.2.
Let $\mathcal{E}_{0}$ be the disjoint collection of intervals of length $2^{-|\mathbf{n}|}$ covering $E_{0}$. Then, we can write
$\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})(x)$ $=\sup_{I\in\mathcal{E}_{0}}\sup_{t\in I}\Bigg|\mathcal{A}_{t}\Big(R_{n_{1}}\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}F_{N}f_{1},R_{n_{2}}f_{2}\Big)(x)\Bigg|,$
where $\widehat{F_{N}(f)}=\chi_{{}_{N}}\widehat{f}$.
For $t\in I$, using the Fundamental Theorem of Calculus we get that
$\sup_{t\in I}\Big|\mathcal{A}_{t}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\Big|$
$\lesssim$ $\left(\frac{1}{|I|}\int_{I}|\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}$
$+\left(\int_{I}\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\frac{\partial}{\partial s}\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\;ds\right)^{\frac{1}{2}}$
$\leq$ $\left(\frac{1}{|I|}\int_{I}|\mathcal{A}_{s}(R_{n_{1}}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}$
$+\left(\int_{I}|\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{4}}\left(\int_{I}\left|\frac{\partial}{\partial s}\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{4}}.$
Using $\ell^{2}\hookrightarrow\ell^{\infty}$ embedding, we get that
$\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})(x)\lesssim$ $\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}2^{\frac{|\mathbf{n}|}{2}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)+\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\big(\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)\big)^{\frac{1}{2}},$ (5.1)
where
$\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}|\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}}~\text{ and}$
$\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|\frac{\partial}{\partial s}\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{2}}.$
Thus, we have
$\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{p}}$
$\lesssim 2^{\frac{|\mathbf{n}|}{2}}\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{p}}+\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{p}}^{\frac{1}{2}}\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{p}}^{\frac{1}{2}}.$ (5.2)
Note that the multiplier $2^{-|\mathbf{n}|}\frac{\partial}{\partial s}\widehat{\sigma}(s\xi,s\eta)$ behaves like $\widehat{\sigma}(s\xi,s\eta)$ where $1\leq s\leq 2$. Therefore, it is enough to prove the desired estimates for the square function $\Big(\sum\limits_{I\in\mathcal{E}_{0}}\int_{I}|\mathcal{A}_{s}(R_{n_{1}}F_{N}f_{1},R_{n_{2}}f_{2})(x)|^{2}\;ds\Big)^{\frac{1}{2}}.$ We invoke the decomposition of the multiplier carried out in Section 4 and use the estimate (4.9) to get that
$\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})(x)$
$\lesssim\left(\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta\right)\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}+error~terms.$
Note that for the error terms, we have better decay in the multiplier, see (4.2). Thus, it is enough to prove $L^{p}$-estimates for the following operator
$f_{1}\rightarrow\left(\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(x,\eta)\right|d\eta\right)$
and
$f_{2}\rightarrow\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}.$
Observe that using Minkowski’s integral inequality, we have
$\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\leq$ $2^{|\mathbf{n}|\frac{\beta}{2}}\int_{{\mathbb{R}}^{d}}\bigg(\int_{I}\left|U_{s,N}^{n_{2}}f_{2}(x-y)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\frac{dy}{\left(1+4\pi^{2}\left|2^{n_{2}}y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}.$ (5.3)
We take $L^{2}$-norm in the above and use Plancherel’s theorem to get that
$\left\|\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{2}}\leq$ $2^{|\mathbf{n}|\frac{\beta}{2}}\left\|\bigg(\int_{I}\big|U_{s,N}^{n_{2}}f_{2}\big|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{2}}\int_{{\mathbb{R}}^{d}}\frac{dy}{\left(1+4\pi^{2}\left|2^{n_{2}}y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}$ (5.4)
$\lesssim$ $2^{-n_{2}d}2^{-|\mathbf{n}|\frac{1-\beta}{2}}\|f_{2}\|_{L^{2}}.$
Further, using the estimate 5.3 along with Lemma 4.2 and Minkowski’s integral inequality, we get that
$\left\|\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\lesssim 2^{-n_{2}\frac{d}{2}}2^{-|\mathbf{n}|\frac{1-\beta}{2}}\|f_{2}\|_{L^{1}}.$ (5.5)
Next, since $\widehat{\sigma}(s\xi,s\eta):=m(s\xi,s\eta)e^{2\pi is|(\xi,\eta)|}$ with $|m(\xi,\eta)|\lesssim|(\xi,\eta)|^{-\frac{2d-1}{2}}$, by an application of Lemma 4.1, we obtain
$\left\|\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(\cdot,\eta)\right|d\eta\right\|_{L^{2}}\lesssim 2^{-\frac{2d-1}{2}|\mathbf{n}|}2^{n_{2}d}\|F_{N}f\|_{L^{2}}.$ (5.6)
Using Hölder’s inequality along with the estimates (5.4), (5.5) and (5.6), we obtain
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{1}}\lesssim$ $2^{-|\mathbf{n}|\frac{2d-\beta}{2}}\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\|F_{N}f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}$
$\lesssim$ $2^{-|\mathbf{n}|\frac{2d-1-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}$
and
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-|\mathbf{n}|\frac{d-\beta}{2}}\bigg(\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\|F_{N}f_{1}\|_{L^{2}}^{\frac{2}{3}}\|f_{2}\|_{L^{1}}^{\frac{2}{3}}\bigg)^{\frac{3}{2}}$
$\lesssim 2^{-|\mathbf{n}|\frac{d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
As described earlier, similar arguments as above, yield the following bounds for square function involving the derivative term $\frac{\partial}{\partial s}\mathcal{A}_{s}$.
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}},$
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-4-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
Substituting the estimates obtained above in (5.1), we have
$\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{1}}\lesssim$ $2^{\frac{|\mathbf{n}|}{2}}2^{-|\mathbf{n}|\frac{2d-1-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}$
$+\left(2^{-|\mathbf{n}|\frac{2d-1-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}\right)^{\frac{1}{2}}\left(2^{-|\mathbf{n}|\frac{2d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}\right)^{\frac{1}{2}}$
$\lesssim$ $2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}},$
and
$\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
Due to symmetry in $\xi$ and $\eta$ variables, we also have
$\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}$ $\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
This completes the proof of Lemma 5.2. ∎
5.2. Proof of Lemma 5.3.
Let $\Psi_{2^{-i}}(w)=2^{id}(2+2^{i}|w|)^{-(d+1)}$ so that $|\psi_{2^{-i}}(w)|\lesssim\Psi_{2^{-i}}(w)$. We note that $\Psi_{2^{-i}}(w_{1})\lesssim\Psi_{2^{-i}}(w_{2})$ whenever $|w_{1}-w_{2}|\leq 2^{-i}$. Indeed, when $|w_{1}-w_{2}|<|w_{2}|$, we have
$\Psi_{2^{-i}}(w_{1})$ $=\Psi_{2^{-i}}(w_{1}-w_{2}+w_{2})$
$=2^{id}(2+2^{i}|w_{1}-w_{2}+w_{2}|)^{-(d+1)}$
$\leq 2^{id}(2+2^{i}(|w_{2}|-|w_{1}-w_{2}|))^{-(d+1)}$
$\leq 2^{id}(1+2^{i}|w_{2}|)^{-(d+1)}=\Psi_{2^{-i}}(w_{2}).$
On the hand hand, when $|w_{2}|\leq|w_{1}-w_{2}|\leq 2^{-i}$, we have
$\Psi_{2^{-i}}(w_{1})$ $\leq 2^{id}(2+2^{i}(|w_{1}-w_{2}|-|w_{2}|))^{-(d+1)}$
$\leq 2^{id}$
$\lesssim 2^{id}(2+2^{i}|w_{2}|)^{-(d+1)}=\Psi_{2^{-i}}(w_{2}).$
Next, we consider the partition ${\mathbb{R}}^{d}=\cup_{Q\in\mathfrak{C}}Q$, where $Q\in\mathfrak{C}$ are cubes with unit sidelengths and sides parallel to co-ordinate axes. For $a<0,$ let $aQ$ denote the cube which is concentric with $Q$ and have sidelength $a>0$. With this we can write
$\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{1/2}}^{\frac{1}{2}}$
$=\int_{{\mathbb{R}}^{d}}\Bigg(\sup\limits_{t\in E_{0}}\bigg|\int_{\mathbb{S}^{2d-1}}\bigg(\sum\limits_{Q\in\mathfrak{C}}(f_{1}*\psi_{2^{-n_{1}}})\chi_{{}_{Q}}\bigg)(x-ty)(f_{2}*\psi_{2^{-n_{2}}})(x-tz)\;d\sigma(y,z)\bigg|\Bigg)^{\frac{1}{2}}dx$
$\leq\sum_{Q\in\mathfrak{C}}\int_{3Q}\bigg(\sup\limits_{t\in E_{0}}\int_{\mathbb{S}^{2d-1}}\big((|f_{1}|*\Psi_{2^{-n_{1}}})\chi_{{}_{Q}}\big)(x-ty)\big((|f_{2}|*\Psi_{2^{-n_{2}}})\chi_{{}_{5Q}}\big)(x-tz)\;d\sigma(y,z)\bigg)^{\frac{1}{2}}dx.$
Let $I\in\mathcal{E}_{0}$ be any interval. Without loss of generality, we may assume that $I=[t_{0}-2^{-|\mathbf{n}|-1},t_{0}+2^{-|\mathbf{n}|-1}]$. For any $t\in I$ and $|y|\leq 1$, we have
$|f_{1}|*\Psi_{2^{-n_{1}}}(x-ty)$ $=|f_{1}|*\Psi_{2^{-n_{1}}}(x-t_{0}y+(t_{0}-t)y)$
$=\int_{{\mathbb{R}}^{d}}|f_{1}|(z)\Psi_{2^{-n_{1}}}(x-t_{0}y+(t_{0}-t)y-z)\;dz$
$\lesssim\int_{{\mathbb{R}}^{d}}|f_{1}|(z)\Psi_{2^{-n_{1}}}(x-t_{0}y-z)\;dz$
$=|f_{1}|*\Psi_{2^{-n_{1}}}(x-t_{0}y),$
where we have used the fact that $\Psi_{2^{-n_{1}}}(w_{1})\lesssim\Psi_{2^{-n_{1}}}(w_{2})$ when $|w_{1}-w_{2}|\leq 2^{-|\mathbf{n}|}\leq 2^{-n_{1}}$ in the last inequality. Similarly, for any $|z|\leq 1$ and $t\in I$, we have
$|f_{2}|*\Psi_{2^{-n_{2}}}(x-tz)\lesssim|f_{2}|*\Psi_{2^{-n_{2}}}(x-t_{0}z).$
Using Hölder’s inequality and the estimates above, we have
$\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\|_{L^{1/2}}^{\frac{1}{2}}$
$\lesssim\sum_{Q\in\mathfrak{C}}\bigg(\int_{3Q}\sum_{I\in\mathcal{E}_{0}}\sup_{t\in I}\int_{\mathbf{S}^{2d-1}}\left((|f_{1}|*\Psi_{2^{-n_{1}}})\chi_{{}_{Q}}\right)(x-ty)\left((|f_{2}|*\Psi_{2^{-n_{2}}})\chi_{{}_{5Q}}\right)(x-tz)\;d\sigma(y,z)\;dx\bigg)^{\frac{1}{2}}$
$\lesssim\sum_{Q\in\mathfrak{C}}\bigg(\int_{3Q}\sum_{I\in\mathcal{E}_{0}}\int_{\mathbf{S}^{2d-1}}\left((|f_{1}|*\Psi_{2^{-n_{1}}})\chi_{{}_{2Q}}\right)(x-t_{0}y)\left((|f_{2}|*\Psi_{2^{-n_{2}}})\chi_{{}_{6Q}}\right)(x-t_{0}z)\;d\sigma(y,z)\;dx\bigg)^{\frac{1}{2}}$
$\lesssim\sum_{Q\in\mathfrak{C}}\bigg(\sum_{I\in\mathcal{E}_{0}}\big\|(|f_{1}|*\Psi_{2^{-n_{1}}})\chi_{{}_{2Q}}\big\|_{L^{1}}\;\big\|(|f_{2}|*\Psi_{2^{-n_{2}}})\chi_{{}_{6Q}}\big\|_{L^{1}}\bigg)^{\frac{1}{2}}$
$\leq 2^{|\mathbf{n}|\frac{\beta}{2}}\bigg(\sum_{Q\in\mathfrak{C}}\|(|f_{1}|*\Psi_{2^{-n_{1}}})\chi_{{}_{2Q}}\|_{L^{1}}\bigg)^{\frac{1}{2}}\bigg(\sum_{Q\in\mathfrak{C}}\|(|f_{2}|*\Psi_{2^{-n_{2}}})\chi_{{}_{6Q}}\|_{L^{1}}\bigg)^{\frac{1}{2}}$
$\lesssim 2^{|\mathbf{n}|\frac{\beta}{2}}\|f_{1}\|_{L^{1}}^{\frac{1}{2}}\|f_{2}\|_{L^{1}}^{\frac{1}{2}},$
where we have used the $L^{1}\times L^{1}\to L^{1}$-boundedness of the single bilinear spherical average $\mathcal{A}_{t}$ obtained in [IPS] in the third step and Cauchy-Schwarz inequality in the fourth step of the equation above. This completes the proof of Lemma 5.3. ∎
6. Estimate for global maximal operator: Proof of Theorem 2.1
Recall that in order to complete the proof of Theorem 2.1 we are required to prove desired $L^{p}$-estimates for the operator
$\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x)=\sup_{k\in{\mathbb{Z}}}\mathcal{M}^{\mathbf{n}}_{k}(f_{1},f_{2})(x).$
Note that since the notion of Minkowski dimension as described in Definition 1.1 is uniform in $k$. Since we are concerned with Hölder exponents $(p_{1},p_{2},p)$, the operator $\mathcal{M}^{\mathbf{n}}_{k}$ satisfies the same bounds as those of $\mathcal{M}^{\mathbf{n}}_{0}$ proved in Lemma 5.2 and Lemma 5.3. Moreover, the bounds are uniform in $k$. Next, we use these estimates for each fixed scale to prove the following results for the operator $\mathcal{M}^{\mathbf{n}}$.
Lemma 6.1.
Let $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
$\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}.$
Lemma 6.2.
Let $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
  1. (1)
    $\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{2/3,\infty}}\lesssim|\mathbf{n}|^{3}2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}},$
  2. (2)
    $\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{2/3,\infty}}\lesssim|\mathbf{n}|^{3}2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}$,
  3. (3)
    $\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\|_{L^{1/2,\infty}}\lesssim|\mathbf{n}|^{4}2^{|\mathbf{n}|\beta}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{1}}$.
Note that if we assume Lemma 6.1 and Lemma 6.2, the proof of Theorem 2.1 follows by arguments similar to those given for the operator $\mathcal{M}_{E_{0}}$.
6.1. Proof of Lemma 6.1.
Recall that $\operatorname{supp}(\widehat{\psi})\subset\big[\frac{1}{2},2\big]$. Let $\boldsymbol{\psi}\in\mathcal{S}({\mathbb{R}}^{d})$ such that $\operatorname{supp}(\widehat{\boldsymbol{\psi}})\subset\big[\frac{1}{4},4\big]$ and $\widehat{\boldsymbol{\psi}}(\xi)=1$ for $\xi\in\big[\frac{1}{2},2\big]$. Define
$\widehat{\mathbf{R}_{i}f}(\xi)=\hat{f}(\xi)\widehat{\boldsymbol{\psi}}_{2^{-i}}(\xi).$
We can rewrite
$\mathcal{M}^{\mathbf{n}}_{k}(f_{1},f_{2})=\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(R_{k+n_{1}}\circ\mathbf{R}_{k+n_{1}}f_{1},R_{k+n_{2}}\circ\mathbf{R}_{k+n_{2}}f_{2})(x)\Big|.$
Using $\ell_{1}\hookrightarrow\ell_{\infty}$ embedding and $L^{2}\times L^{2}\to L^{1}$-estimate from Lemma 5.2, we have
$\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{1}}$ $\leq\left\|\sum_{k\in{\mathbb{Z}}}\mathcal{M}^{\mathbf{n}}_{k}(f_{1},f_{2})\right\|_{L^{1}}$
$\lesssim\sum_{k\in{\mathbb{Z}}}2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|\mathbf{R}_{k+n_{1}}f_{1}\|_{L^{2}}\|\mathbf{R}_{k+n_{2}}f_{2}\|_{L^{2}}$
$\leq 2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\left(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{1}}f_{1}\|_{L^{2}}^{2}\right)^{\frac{1}{2}}\left(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{2}}f_{2}\|_{L^{2}}^{2}\right)^{\frac{1}{2}}$
$\lesssim 2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}.$
This completes the proof of Lemma 6.1. ∎
6.2. Proof of parts $(1)$ and $(2)$ of Lemma 6.2
Note that due to symmetry it is enough to prove part $(1)$. The proof is based on a suitable adaptation of the bilinear Calderón-Zygmund theory and delicate analysis involving the exceptional set which arises in the Calderón-Zygmund decomposition of input function. Without loss of generality we may assume that $\|f_{1}\|_{L^{1}}=\|f_{2}\|_{L^{2}}=1$. Given $\alpha>0$, we shall prove the following
$\Big|\{x:\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x)>\alpha\}\Big|\leq C|\mathbf{n}|^{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{-\frac{2}{3}}.$ (6.1)
We apply the Calderón-Zygmund decomposition to $f_{1}$ at height $2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{\frac{2}{3}}$ to obtain the decomposition $f_{1}=g_{1}+h_{1}$ with a collection $\{Q_{\gamma}\}$ of disjoint cubes such that
$\|g_{1}\|_{L^{\infty}}\leq 2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{\frac{2}{3}},~~h_{1}=\sum_{\gamma}h_{\gamma},~~supp(h_{\gamma})\subset Q_{\gamma},$
$\|h_{\gamma}\|_{L^{1}}\lesssim 2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{\frac{2}{3}}|Q_{\gamma}|~~\text{with}~~\int_{Q_{\gamma}}h_{\gamma}=0,~~\text{and}$
$\sum_{\gamma}\big|Q_{\gamma}\big|\leq 2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{-\frac{2}{3}}.$
Note that we have
$\Big|\{x:\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x)>\alpha\}\Big|\leq$ $\Big|\{x:\mathcal{M}^{\mathbf{n}}(h_{1},f_{2})(x)>\frac{\alpha}{2}\}\Big|+\Big|\{x:\mathcal{M}^{\mathbf{n}}(g_{1},f_{2})(x)>\frac{\alpha}{2}\}\Big|.$
Contribution from $\mathcal{M}^{\mathbf{n}}(g_{1},f_{2})$:
This is the easy part. We use Chebyshev’s inequality and Lemma 6.1 to see that
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},f_{2})(x)>\frac{\alpha}{2}\}|\lesssim\frac{2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|g_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}}{\alpha}.$
Since $\|g_{1}\|_{2}\lesssim 2^{|\mathbf{n}|\frac{d-3-\beta}{6}}\alpha^{\frac{1}{3}}$, we obtain
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},f_{2})(x)>\frac{\alpha}{2}\}|$ $\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{-\frac{2}{3}}.$
Contribution from $\mathcal{M}^{\mathbf{n}}(h_{1},f_{2})$:
Let $\tilde{Q}_{\gamma_{j}}$ denote the cube which is concentric with ${Q}_{\gamma_{j}}$ and have sidelength $l(\tilde{Q}_{\gamma_{j}})=10l(Q_{\gamma_{j}})$. We define the exceptional set
$\mathcal{F}:=\Big(\cup_{\gamma}\tilde{Q}_{\gamma}\Big).$
Then, $|\mathcal{F}|\leq 2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{-\frac{2}{3}}$. Thus, by Chebyshev’s inequality, it is enough to show that
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}|\mathcal{M}^{\mathbf{n}}(h_{1},f_{2})(x)|^{\frac{2}{3}}~dx\lesssim|\mathbf{n}|^{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}.$ (6.2)
We organize the function $h_{1}$ as follows
$h_{1}=\sum_{i\in\mathbb{Z}}h^{i}_{1},~~\text{where}~~h^{i}_{1}=\sum_{l(Q_{\gamma})=2^{-i}}h_{\gamma}.$
Lemma 6.3.
Let $k,i\in\mathbb{Z}$ and $\mathbf{n}\in\mathbb{N}^{2}$. Then the following estimate holds uniformly in $k,i$ and $\mathbf{n}$.
$\int\limits_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx\lesssim n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\min\big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\big\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
We assume Lemma 6.3 for a moment and complete the proof. In fact, we can obtain the following estimate in Lemma 6.3.
$\int\limits_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx\lesssim n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\min\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
Observe that
$\sum_{i\in{\mathbb{Z}}}\min\big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\big\}=\sum_{k\in{\mathbb{Z}}}\min\big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\big\}\lesssim n_{1}.$
Using the embedding $\ell^{1}\hookrightarrow\ell^{\infty}$ and the above estimates, we have
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}|\mathcal{M}^{\mathbf{n}}(h_{1},f_{2})(x)|^{\frac{2}{3}}dx$
$\leq\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\left|\sum_{k\in{\mathbb{Z}}}\mathcal{M}^{\mathbf{n}}_{k}\Big(\sum_{i\in{\mathbb{Z}}}h_{1}^{i},f_{2}\Big)(x)\right|^{\frac{2}{3}}dx$
$\leq\sum_{k,i\in{\mathbb{Z}}}\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\left|\mathcal{M}^{\mathbf{n}}_{k}\big(h_{1}^{i},f_{2}\big)(x)\right|^{\frac{2}{3}}dx$
$\lesssim\sum_{k,i\in{\mathbb{Z}}}n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}$
$\lesssim n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\sum_{k\in{\mathbb{Z}}}\bigg(\sum_{i\in{\mathbb{Z}}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\bigg)^{\frac{1}{3}}$
$\hskip 91.04872pt\times\bigg(\sum_{i\in{\mathbb{Z}}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\|h^{i}_{1}\|_{L^{1}}\bigg)^{\frac{2}{3}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}$
$\leq n_{2}n_{1}^{\frac{1}{3}}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\bigg(\sum_{k\in{\mathbb{Z}}}\sum_{i\in{\mathbb{Z}}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\|h^{i}_{1}\|_{L^{1}}\bigg)^{\frac{2}{3}}\bigg(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{2}_{L^{2}}\bigg)^{\frac{1}{3}}$
$\lesssim n_{2}n_{1}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\bigg(\sum_{i\in{\mathbb{Z}}}\|h^{i}_{1}\|_{L^{1}}\bigg)^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\lesssim|\mathbf{n}|^{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}},$
where we have used Hölder’s inequality in the fourth and fifth steps with respect to $i$ and $k$ respectively.
This completes the proof of part $(1)$ of Lemma 6.2 under the assumption that Lemma 6.3 holds.
Proof of Lemma 6.3.
We apply part $(2)$ of Lemma 5.2 to get that
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx$ $\lesssim$ $2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\|\mathbf{R}_{k+n_{1}}h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}$
$\lesssim$ $2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
Observe that this gives us the desired estimate for the terms satisfying $k\leq i<k+n_{1}$.
Next, we consider terms with $i\geq k+n_{1}$. We use the cancellation condition of $h_{\gamma}$ to write
$\mathbf{R}_{k+n_{1}}h_{\gamma}(x)$ $=\int_{Q_{\gamma}}\Big(\boldsymbol{\psi}_{2^{-(k+n_{1})}}(x-y)-\boldsymbol{\psi}_{2^{-(k+n_{1})}}(x-c_{Q})\Big)h_{\gamma}(y)dy$
$=\int_{\mathbb{R}^{d}}\int^{1}_{0}\langle\nabla_{y}\boldsymbol{\psi}_{2^{-(k+n_{1})}}(x-c_{Q}-t(y-c_{Q})),y-c_{Q}\rangle dt~h_{\gamma}(y)dy,$
where $c_{Q}$ is the center of $Q_{\gamma}$. We observe that for $y\in Q_{\gamma}$, we have
$\Big|\langle\nabla_{y}\boldsymbol{\psi}_{2^{-(k+n_{1})}}(x-c_{Q}-t(y-c_{Q})),y-c_{Q}\rangle\Big|\lesssim{2^{(k+n_{1}-i)}}|(\nabla_{y}\boldsymbol{\psi})_{2^{-(k+n_{1})}}(x-c_{Q}-t(y-c_{Q}))|.$
Applying Minkowski’s integral inequality, we get
$\|\mathbf{R}_{k+n_{1}}h^{i}_{1}\|_{L^{1}}\lesssim 2^{k+n_{1}-i}\|h^{i}_{1}\|_{L^{1}}.$
Finally, for the remaining case when $i<k$, we exploit the restriction of the domain of integration to the complement of $\mathcal{F}$. Recall that $h^{i}_{1}=\sum\limits_{l(Q_{\gamma})=2^{-i}}h_{\gamma}$. Then, for $y^{\prime}\in\operatorname{supp}(h_{\gamma})$ and $x\notin\mathcal{F}$, we have
$dist(x-ty,y^{\prime})\geq 2^{2-i}\quad\text{for all }t\in E_{k},~|y|\leq 1,$
as $2^{-k}<2^{-i}$. Consequently, we have
$\|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})\|_{L^{2/3}(\mathbb{R}^{d}\setminus\mathcal{F})}^{\frac{2}{3}}$
$=\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\sup_{t\in E_{k}}\Big|\int_{\mathbb{S}^{2d-1}}h_{1}^{i}\ast\psi_{2^{-k-n_{1}}}(x-ty)f_{2}\ast\psi_{2^{-k-n_{2}}}(x-tz)~d\sigma(y,z)\Big|^{\frac{2}{3}}dx$
$=\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\sup_{t\in E_{k}}\bigg|\int_{\mathbb{S}^{2d-1}}\Big(\int_{{\mathbb{R}}^{d}}h_{1}^{i}(y^{\prime})\psi_{2^{-k-n_{1}}}(x-ty-y^{\prime})~dy^{\prime}\Big)f_{2}\ast\psi_{2^{-k-n_{2}}}(x-tz)~d\sigma(y,z)\bigg|^{\frac{2}{3}}dx$
$=\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\sup_{t\in E_{k}}\bigg|\int_{\mathbb{S}^{2d-1}}\Big(\int_{{\mathbb{R}}^{d}}h_{1}^{i}(y^{\prime})\big(\psi_{2^{-k-n_{1}}}\chi_{\{|\cdot|\geq 2^{-i}\}}\big)(x-ty-y^{\prime})~dy^{\prime}\Big)f_{2}\ast\psi_{2^{-k-n_{2}}}(x-tz)~d\sigma(y,z)\bigg|^{\frac{2}{3}}dx$
$\leq\bigg\|\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},R_{k+n_{2}}f_{2})\Big|\bigg\|_{L^{2/3}(\mathbb{R}^{d})}^{\frac{2}{3}},$
where $\varphi_{2^{-k-n_{1}}}^{i}(w)=\psi_{2^{-k-n_{1}}}(w)\chi_{\{|w|\geq 2^{-i}\}}(w)$.
Using partition of unity from Section 3, we can decompose
$\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},R_{k+n_{2}}f_{2})(x)\Big|\leq\sum_{m\geq 0}\mathcal{M}^{(m,n_{2})}_{k}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},f_{2})(x),$
where
$\mathcal{M}^{(m,n_{2})}_{k}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},f_{2})(x)=\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}\big(R_{k+m}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}),R_{k+n_{2}}f_{2}\big)(x)\Big|$
for $m\geq 1$ and
$\mathcal{M}^{(0,n_{2})}_{k}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},f_{2})(x)=\sup_{t\in E_{k}}\Big|\mathcal{A}_{t}\big(P_{k}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}),R_{k+n_{2}}f_{2}\big)(x)\Big|.$
For $m\geq 0$, we claim that
$\|\mathcal{M}^{(m,n_{2})}_{k}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},f_{2})\|_{L^{2/3}}\lesssim 2^{-|(m,n_{2})|\frac{d-3-\beta}{2}}\|h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$ (6.3)
Let us assume the estimate (6.3) for a moment. Note that $\psi\in\mathcal{S}({\mathbb{R}}^{d})$, we have $|\psi(w)|\lesssim_{N}(1+|w|)^{-N}$ for every $N>1$. Consequently, for $N\geq d+1$, we get that
$\|\varphi_{2^{-k-n_{1}}}^{i}\|_{L^{1}}$ $=\int_{|w|\geq 2^{-i}}\psi_{2^{-k-n_{1}}}(w)\;dw$
$\lesssim\int_{|w|\geq 2^{-i}}2^{(k+n_{1})d}(2+2^{k+n_{1}}|w|)^{-N}\;dw$
$=\int_{|w|\geq 2^{-i+k+n_{1}}}(2+|w|)^{-N}\;dw$
$\lesssim 2^{(N-d)(i-n_{1}-k)}$
$\leq 2^{-(N-d)n_{1}}2^{i-k},$
where the last inequality follows from the condition $i<k$. We choose $N=2d$ to obtain
$\|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})\|_{L^{2/3}(\mathbb{R}^{d}\setminus\mathcal{F})}^{\frac{2}{3}}$ $\leq\sum_{m\geq 0}\|\mathcal{M}^{(m,n_{2})}_{k}(h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i},f_{2})\|_{L^{2/3}({\mathbb{R}}^{d})}^{\frac{2}{3}}$
$\lesssim\sum_{m\geq 0}2^{-|(m,n_{2})|\frac{d-3-\beta}{3}}\|h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\lesssim\sum_{m\geq 0}2^{-|(m,n_{2})|\frac{d-3-\beta}{3}}2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\leq 2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}\bigg(\sum_{m=0}^{n_{2}-1}2^{-n_{2}\frac{d-3-\beta}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}+\sum_{m\geq n_{2}}2^{-m\frac{d-3-\beta}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}\bigg)$
$\leq n_{2}2^{-\frac{2n_{1}d}{3}}2^{\frac{2(i-k)}{3}}2^{-n_{2}\frac{d-3-\beta}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}$
$\leq n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}2^{\frac{2(i-k)}{3}}\|h^{i}_{1}\|_{L^{1}}^{\frac{2}{3}}\|f_{2}\|_{L^{2}}^{\frac{2}{3}}.$
This completes the proof of Lemma 6.3 given that the estimate (6.3) holds.
Proof of estimate (6.3):
When $m\geq 1$, the desired estimate in (6.3) follows by adapting the proof of Lemma 5.2 to the scale $k$. Thus, it remains to prove (6.3) for $m=0$.
Let $b_{1}=h^{i}_{1}\ast\varphi_{2^{-k-n_{1}}}^{i}$. In view of the discussion in Section 4, we need to prove the desired $L^{p}$-boundedness for the maximal operator corresponding to bilinear operators defined by
$T^{n_{2}}_{t,k}(b_{1},f_{2})(x)=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{P_{k}b_{1}}(\xi)\widehat{R_{k+n_{2}}f_{2}}(\eta)m(t\xi,t\eta)e^{2\pi it|(\xi,\eta)|}e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,\;t\in E_{k},$
where $m(\xi,\eta)=|(\xi,\eta)|^{-\frac{2d-1}{2}}$. Again, we restrict ourselves to the case of unit scale, that is $k=0$. The following arguments can be easily adapted to obtain the corresponding estimates for a general scale $k$. As previously, we denote $T^{n_{2}}_{t,0}$ by $T^{n_{2}}_{t}$.
Let $N$ be a positive integer such that $N\sim 2^{n_{2}}$ and write
$e^{2\pi it|(\xi,\eta)|}=e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi it\sqrt{|\xi|^{2}+N^{2}}},$
where
$\Phi_{N}(\xi,\eta):=|(\xi,\eta)|-\sqrt{|\xi|^{2}+N^{2}}=\frac{(|\eta|-N)(|\eta|+N)}{\sqrt{|\xi|^{2}+|\eta|^{2}}+\sqrt{|\xi|^{2}+N^{2}}}.$
Note that $\Phi_{N}$ satisfies the following estimate
$|\partial^{\alpha}_{\xi}\Phi_{N}(\xi,\eta)|\lesssim 1,$ (6.4)
for all multi-indices $\alpha$ uniformly in $(\xi,\eta)$ when $|\xi|\leq 2$ and $N\leq|\eta|<N+1$. This gives us the following decomposition of the operator $T^{n_{2}}_{t}.$
$T^{n_{2}}_{t}(b_{1},f_{2})(x)$ $=\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}T^{n_{2}}_{t,N}(b_{1},f_{2})(x),$
where
$T^{n_{2}}_{t,N}(b_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}m_{t}^{n_{2}}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\widehat{U_{t,N}b_{1}}(\xi)\widehat{F_{N}f_{2}}(\eta)e^{2\pi ix\cdot(\xi,\eta)}\;d\xi d\eta,$
$\widehat{U_{t,N}b_{1}}(\xi):=\widehat{b_{1}}(\xi)\widehat{\phi}\left(\xi\right)e^{2\pi it\left(\sqrt{|\xi|^{2}+N^{2}}\right)},$ (6.5)
and
$m^{n_{2}}_{t}(\xi,\eta):=m(t\xi,t\eta)\widehat{\phi}\left(\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right).$
Using the estimate (6.4), we get that
$|\partial^{\alpha}_{\xi}\big(m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\big)|\lesssim C_{\alpha}2^{-\frac{2d-1}{2}|n_{2}|},$ (6.6)
for all multi-indices $\alpha$ uniformly in $t\in E_{0}$.
Since $\widehat{U_{t,N}b_{1}}(\eta)=\int_{{\mathbb{R}}^{d}}U_{t,N}b_{1}(y)e^{-2\pi i\xi\cdot y}~dy$, we can rewrite
$T^{n_{2}}_{t,N}(b_{1},f_{2})(x)$ $=\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{F_{N}f_{2}}(\eta)U_{t,N}b_{1}(y)\Big(\int_{{\mathbb{R}}^{d}}m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi i\xi\cdot(x-y)}d\xi\Big)e^{2\pi ix\cdot\eta}\;d\eta dy.$ (6.7)
Similarly to (4.7), an integration by parts argument gives
$\int_{{\mathbb{R}}^{d}}m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}e^{2\pi i\xi\cdot(x-y)}d\xi$ (6.8)
$=\frac{1}{(1+4\pi^{2}|x-y|^{2})^{\lceil\frac{d}{2}\rceil+1}}\int_{{\mathbb{R}}^{d}}(1-\Delta_{\xi})^{\lceil\frac{d}{2}\rceil+1}\Big(m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\Big)e^{2\pi i\xi\cdot(x-y)}d\xi,$
Substituting (6.8) into (6.7), we get
$T^{n_{2}}_{t,N}(b_{1},f_{2})(x)$ $=\int_{\mathbb{R}^{d}}A\left(U_{t,N}b_{1}\right)(x,\xi)B^{n_{2}}_{t,N}\left(F_{N}f_{2}\right)(x,\xi)d\xi,$
where
$B^{n_{2}}_{t,N}\left(F_{N}f_{2}\right)(x,\xi):=$ $\int_{\mathbb{R}^{d}}\left(1-\Delta_{\xi}\right)^{\lceil\frac{d}{2}\rceil+1}\left(m^{n_{2}}_{t}(\xi,\eta)e^{2\pi it\Phi_{N}(\xi,\eta)}\right)\widehat{F_{N}f_{2}}(\eta)e^{2\pi ix\cdot\eta}d\eta,$ (6.9)
and
$A\left(U_{t,N}b_{1}\right)(x,\eta):=\int_{\mathbb{R}^{d}}\frac{U_{t,N}b_{1}(y)e^{2\pi i\xi\cdot(x-y)}}{\left(1+4\pi^{2}\left|x-y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy.$
Note that
$|A\left(U_{t,N}b_{1}\right)(x,\eta)|$ $\leq\int_{\mathbb{R}^{d}}\frac{|U_{t,N}b_{1}(y)|}{\left(1+4\pi^{2}\left|x-y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
$:=L\left(U_{t,N}b_{1}\right)(x).$
This gives us the following bound.
$|T^{n_{2}}_{t,N}(b_{1},f_{2})(x)|\leq L\left(U_{t,N}b_{1}\right)(x)\int_{\mathbb{R}^{d}}\left|B^{n_{2}}_{t,N}\left(F_{N}f_{2}\right)(x,\xi)\right|d\xi.$ (6.10)
With this decomposition in mind, we follow the arguments from the proof of estimate (5.1) to get that
$\mathcal{M}^{(0,n_{2})}_{0}(b_{1},f_{2})(x)\lesssim$ $\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}2^{\frac{n_{2}}{2}}\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})(x)+\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\big(\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})(x)\tilde{\mathcal{G}}^{n_{2}}_{0,N}(b_{1},f_{2})(x)\big)^{\frac{1}{2}},$
where
$\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}|\mathcal{A}_{s}(P_{0}b_{1},R_{n_{2}}F_{N}f_{2})(x)|^{2}\;ds\right)^{\frac{1}{2}},$
$\tilde{\mathcal{G}}^{n_{2}}_{0,N}(b_{1},f_{2})(x):=\left(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|\frac{\partial}{\partial s}\mathcal{A}_{s}(P_{0}b_{1},R_{n_{2}}F_{N}f_{2})(x)\right|^{2}\;ds\right)^{\frac{1}{2}},$
and $\mathcal{E}_{0}$ is a collection of disjoint intervals of length $2^{-n_{2}}$ and covering $E_{0}$.
Since the multiplier $2^{-|\mathbf{n}|}\frac{\partial}{\partial s}\widehat{\sigma}(s\xi,s\eta)$ behaves like $\widehat{\sigma}(s\xi,s\eta)$ where $1\leq s\leq 2$. Therefore, it is enough to prove the desired estimates for the square function $\Big(\sum\limits_{I\in\mathcal{E}_{0}}\int_{I}|\mathcal{A}_{s}(P_{0}f_{1},R_{n_{2}}F_{N}f_{2})(x)|^{2}\;ds\Big)^{\frac{1}{2}}$.
We invoke the decomposition of the multiplier in (6.10) to get that
$\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})(x)$
$\lesssim\left(\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{n_{2}}_{s,N}\left(F_{N}f_{2}\right)(x,\xi)\right|d\xi\right)\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L\left(U_{s,N}b_{1}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}+error~terms.$
Note that for the error terms, we have better decay of the multiplier, see (4.2). Thus, it is enough to prove desired bounds for the leading term.
Recall that $\operatorname{supp}(\widehat{\phi})\subset B(0,2)$. Similar to the $L^{p}$-estimates in Lemma 4.1, we can get that
$\left\|\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{n_{2}}_{s,N}\left(F_{N}f_{2}\right)(\cdot,\xi)\right|d\xi\right\|_{L^{2}}\lesssim 2^{-n_{2}\frac{2d-1}{2}}\|f_{2}\|_{L^{2}}.$
Using Minkowski’s integral inequality twice, we have
$\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L\left(U_{s,N}b_{1}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}$
$\leq 2^{n_{2}\frac{\beta}{2}}\int_{{\mathbb{R}}^{d}}\bigg(\int_{I}\left|U_{s,N}b_{1}(x-y)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\frac{1}{\left(1+4\pi^{2}\left|y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}dy$
$\leq 2^{n_{2}\frac{\beta}{2}}\int_{{\mathbb{R}}^{d}}\int_{{\mathbb{R}}^{d}}b_{1}(x-y-z)\Big(\int_{I}\left|K_{s,N}(z)\right|^{2}\;ds\Big)^{\frac{1}{2}}\frac{1}{\left(1+4\pi^{2}\left|y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}~dzdy,$
where $K_{s,N}(z)=\int_{{\mathbb{R}}^{d}}\widehat{\phi}\left(\xi\right)e^{2\pi is\left(\sqrt{|\xi|^{2}+N^{2}}\right)}e^{2\pi iz\cdot\xi}d\xi$. Since $\phi\in\mathcal{S}({\mathbb{R}}^{d})$ and $s\in I\subset[1,2]$, we have good kernel estimates for $K_{s,N}$ which are uniform in $s$. Thus, we have
$\left\|\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L\left(U_{s,N}b_{1}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\lesssim 2^{-n_{2}\frac{1-\beta}{2}}\|b_{1}\|_{L^{1}}.$
Using Hölder’s inequality and the estimates above, we get that
$\Bigg\|\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\mathcal{G}^{n_{2}}_{0,N}(b_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-n_{2}|\frac{2d-\beta}{2}}\bigg(\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\|b_{1}\|_{L^{1}}^{\frac{2}{3}}\|F_{N}f_{2}\|_{L^{2}}^{\frac{2}{3}}\bigg)^{\frac{3}{2}}$
$\lesssim 2^{-n_{2}\frac{2d-\beta-2}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
Similarly, we can get
$\Bigg\|\sum_{N=2^{n_{2}-1}}^{2^{n_{2}+1}}\tilde{\mathcal{G}}^{n_{2}}_{0,N}(b_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-n_{2}\frac{2d-\beta-4}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
Consequently, we have
$\big\|\mathcal{M}^{(0,n_{2})}_{0}(b_{1},f_{2})\big\|_{L^{2/3}}\lesssim 2^{-n_{2}\frac{2d-\beta-3}{2}}\|b_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
This proves the desired bounds in (6.3) for $\mathcal{M}^{(0,n_{2})}_{0}$. The above argument can be modified suitably to get the desired bounds for $\mathcal{M}^{(0,n_{2})}_{k}$. ∎
Proof of part $(3)$ of Lemma 6.2.
Let us assume without loss of generality that $\|f_{1}\|_{L^{1}}=\|f_{2}\|_{L^{1}}=1$. Hence, for given $\alpha>0$, we are required to prove that
$\Big|\{x:\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x)>\alpha\}\Big|\leq C|\mathbf{n}|^{2}2^{|\mathbf{n}|\frac{\beta}{2}}\alpha^{-\frac{1}{2}}.$ (6.11)
We apply the Calderón-Zygmund decomposition to both functions $f_{1}$ and $f_{2}$ with the same choice of height $\alpha^{\frac{1}{2}}$. This gives us the decomposition $f_{j}=g_{j}+h_{j}$ with collections $\{Q_{\gamma_{j}}\}$ of disjoint cubes for $j=1,2$ such that
$\|g_{j}\|_{L^{\infty}}\leq\alpha^{\frac{1}{2}},~~h_{j}=\sum_{\gamma_{j}}h_{\gamma_{j}},~~supp(h_{\gamma_{j}})\subset Q_{\gamma_{j}},$
$\|h_{\gamma_{j}}\|_{L^{1}}\leq C\alpha^{\frac{2}{3}}|Q_{\gamma_{j}}|~~\text{with}~~\int_{Q_{\gamma_{j}}}h_{\gamma_{j}}=0,~~\text{and}~~\sum_{\gamma_{j}}\big|Q_{\gamma_{j}}\big|\leq\alpha^{-\frac{1}{2}}.$
We have
$\Big|\{x:\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x)>\alpha\}\Big|\leq$ $\Big|\{x:\mathcal{M}^{\mathbf{n}}(h_{1},h_{2})(x)>\frac{\alpha}{4}\}\Big|+\Big|\{x:\mathcal{M}^{\mathbf{n}}(h_{1},g_{2})(x)>\frac{\alpha}{4}\}\Big|$
$+\Big|\{x:\mathcal{M}^{\mathbf{n}}(g_{1},h_{2})(x)>\frac{\alpha}{4}\}\Big|+\Big|\{x:\mathcal{M}^{\mathbf{n}}(g_{1},g_{2})(x)>\frac{\alpha}{4}\}\Big|.$
We need to prove the estimate in (6.11) for each of the four terms in the equation above.
Contribution from $\mathcal{M}^{\mathbf{n}}(g_{1},g_{2})$:
From Chebyshev’s inequality and Lemma 6.1, we get that
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},g_{2})(x)>\frac{\alpha}{4}\}|$ $\lesssim$ $\frac{2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|g_{1}\|_{L^{2}}\|g_{2}\|_{L^{2}}\|_{L^{1}}}{\alpha}$
$\lesssim$ $\alpha^{-\frac{1}{2}},$
where we used $\|g_{1}\|_{L^{2}}\lesssim\alpha^{\frac{1}{4}}$ and $\|g_{2}\|_{L^{2}}\lesssim\alpha^{\frac{1}{4}}$ in the last step. Thus, we obtain (6.11) for $\mathcal{M}^{\mathbf{n}}(g_{1},g_{2})$.
Contributions from $\mathcal{M}^{\mathbf{n}}(h_{1},g_{2})$ and $\mathcal{M}^{\mathbf{n}}(g_{1},h_{2})$:
Due to symmetry it is enough to provide the arguments for the term $\mathcal{M}^{\mathbf{n}}(h_{1},g_{2})$. Note that by part $(1)$ of Lemma 6.2, we get
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(h_{1},g_{2})(x)>\frac{\alpha}{4}\}|$ $\lesssim$ $\frac{|\mathbf{n}|^{2}\|h_{1}\|_{L^{1}}^{\frac{2}{3}}\|g_{2}\|_{L^{2}}^{\frac{2}{3}}}{\alpha^{\frac{2}{3}}}$
$\lesssim$ $\frac{|\mathbf{n}|^{2}\alpha^{\frac{1}{6}}}{\alpha^{\frac{2}{3}}}$
$=$ $|\mathbf{n}|^{2}\alpha^{-\frac{1}{2}},$
where we have used the fact $\|g_{2}\|_{L^{2}}\lesssim\alpha^{\frac{1}{4}}$ in the second inequality. This proves the desired estimates for $\mathcal{M}^{\mathbf{n}}(h_{1},g_{2})$.
Contribution from $\mathcal{M}^{\mathbf{n}}(h_{1},h_{2})$:
We use the notation as in the previous section. We define the exceptional set
$\mathcal{F}:=\Big(\cup_{\gamma_{1}}\tilde{Q}_{\gamma_{1}}\Big)\cup\Big(\cup_{\gamma_{2}}\tilde{Q}_{\gamma_{2}}\Big).$
Then, $|\mathcal{F}|\lesssim\alpha^{-\frac{1}{2}}$. Thus, using Chebyshev’s inequality, it is enough to show that
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}|\mathcal{M}^{\mathbf{n}}(h_{1},h_{2})(x)|^{\frac{1}{2}}~dx\lesssim|\mathbf{n}|^{2}2^{|\mathbf{n}|\frac{\beta}{2}}.$
We organize the bad functions
$h_{j}=\sum_{i_{j}\in\mathbb{Z}}h^{i_{j}}_{j},~~\text{where}~~h^{i_{j}}_{j}=\sum_{l(Q_{\gamma_{j}})=2^{-i}}h_{\gamma_{j}}.$
Lemma 6.4.
Let $k,i_{1},i_{2}\in\mathbb{Z}$ and $\mathbf{n}\in\mathbb{N}^{2}$. Then the following estimate holds uniformly in $k,i_{j},n_{j}$.
$\int\limits_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i_{1}}_{1},h^{i_{2}}_{2})(x)\Big|^{\frac{1}{2}}dx\lesssim\min_{j=1,2}\{1,2^{\frac{i_{j}-k}{2}},2^{\frac{k+n_{j}-i_{j}}{2}}\}2^{|\mathbf{n}|\frac{\beta}{2}}\prod^{2}_{j=1}\|h^{i_{j}}_{j}\|^{\frac{1}{2}}_{L^{1}}.$
We use similar arguments as in the previous section along with Lemma 6.4 and Lemma $5.2$ from [ChristZhou] to conclude that
$\sum_{k\in\mathbb{Z}}\sum_{i_{1},i_{2}\in\mathbb{Z}}\min_{j=1,2}\min\Big\{2^{\frac{i_{j}-k}{2}},2^{\frac{k+n_{j}-i_{j}}{2}},1\Big\}\prod^{2}_{j=1}\|h^{i_{j}}_{j}\|^{\frac{1}{2}}_{L^{1}}\lesssim|\mathbf{n}|^{2}2^{|\mathbf{n}|\frac{\beta}{2}}.$
This completes the proof for the parts where we have bad functions in both places.
Proof of Lemma 6.4.
The proof is similar to that of Lemma 6.3 with some modifications. We will give a brief sketch only. Note that from Lemma 5.3, we have
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i_{1}}_{1},h^{i_{2}}_{2})(x)\Big|^{\frac{1}{2}}dx$ $\lesssim$ $2^{|\mathbf{n}|\frac{\beta}{2}}\|h^{i_{1}}_{1}\|^{\frac{1}{2}}_{L^{1}}\|h^{i_{2}}_{2}\|^{\frac{1}{2}}_{L^{1}}.$
The estimate above gives us the desired bound when $k\leq i_{j}<k+n_{j}$.
When $i_{j}>k+n_{j}$, we write $h^{i_{j}}_{j}=\sum\limits_{l(Q_{\gamma_{j}})=2^{-i_{j}}}h_{\gamma_{j}}$. Using the mean zero property of the functions $h_{\gamma_{j}}$, we get
$\|\mathbf{R}_{k+n_{j}}h^{i_{j}}_{j}\|_{L^{1}}\lesssim 2^{k+n_{j}-i_{j}}\|h^{i_{j}}_{j}\|_{L^{1}}.$ (6.12)
When $i_{j}<k$, we follow the idea as used in the proof of Lemma 5.3 with some modifications. Write ${\mathbb{R}}^{d}=\cup_{Q\in\mathfrak{C}}Q$, where $\mathfrak{C}$ is a family of disjoint cubes with sidelength $2^{-k}$ and having sides parallel to the coordinate axes. Then
$\|\mathcal{M}^{\mathbf{n}}_{k}(h^{i_{1}}_{1},h^{i_{2}}_{2})\|_{L^{1/2}(\mathbb{R}^{d}\setminus\mathcal{F})}^{\frac{1}{2}}$
$=\int_{\mathbb{R}^{d}\setminus\mathcal{F}}\bigg(\sup\limits_{t\in E_{k}}\bigg|\int_{\mathbb{S}^{2d-1}}\bigg(\sum\limits_{Q\in\mathfrak{C}}(h^{i}_{1}*\psi_{2^{-k-n_{1}}})\chi_{{}_{Q}}\bigg)(x-ty)(f_{2}*\psi_{2^{-k-n_{2}}})(x-tz)\;d\sigma(y,z)\bigg|\bigg)^{\frac{1}{2}}dx$
$\leq\sum_{Q\in\mathfrak{C}}\int_{3Q\setminus\mathcal{F}}\bigg(\sup\limits_{t\in E_{k}}\int_{\mathbb{S}^{2d-1}}\big((|h^{i_{1}}_{1}|*\Psi_{2^{-k-n_{1}}})\chi_{{}_{Q}}\big)(x-ty)\big((|h^{i_{2}}_{2}|*\Psi_{2^{-k-n_{2}}})\chi_{{}_{5Q}}\big)(x-tz)\;d\sigma(y,z)\bigg)^{\frac{1}{2}}dx.$
where $\Psi(w)=(2+|w|)^{-N}$ so that $|\psi(w)|\lesssim\Psi(w)$. Recall that $h^{i_{1}}_{1}=\sum\limits_{l(Q_{\gamma_{1}})=2^{-i_{1}}}h_{\gamma_{1}}$. Then, for $y^{\prime}\in\operatorname{supp}(h_{\gamma_{1}})$ and $x\notin\mathcal{F}$, we have
$dist(x-ty,y^{\prime})\geq 2^{2-i_{1}}\quad\text{for all }t\in E_{k},~|y|\leq 1,$
as $2^{-k}<2^{-i_{1}}$.
We define $\Phi_{2^{-k-n_{1}}}(w)=\Psi_{2^{-k-n_{1}}}(w)\chi_{\{|w|\geq 2^{-i_{1}}\}}(w)$. Therefore, when $x\notin\mathcal{F}$, we have
$|h^{i_{1}}_{1}|*\Psi_{2^{-k-n_{1}}}(x-ty)\lesssim|h^{i_{1}}_{1}|*\Phi_{2^{-k-n_{1}}}(x-ty)\quad\text{for all }t\in E_{k},~|y|\leq 1.$
Recall that for any $N>d$, we have
$\|\Phi_{2^{-n_{1}-k}}\|_{L^{1}}$ $\lesssim 2^{(N-d)(i_{1}-n_{1}-k)}.$
Let $\mathcal{E}_{k}$ be the collection of disjoint intervals of length $2^{-k-|\mathbf{n}|}$ covering $E_{k}$. Recall that
$\Psi_{2^{-k-n_{1}}}(w_{1})\lesssim\Psi_{2^{-k-n_{1}}}(w_{2}),~\text{ whenever}~|w_{1}-w_{2}|\leq 2^{-k-|\mathbf{n}|}.$
Now, using the Cauchy-Schwarz inequality, we have
$\|\mathcal{M}^{\mathbf{n}}_{k}(h^{i_{1}}_{1},h^{i_{2}}_{2})\|_{L^{1/2}(\mathbb{R}^{d}\setminus\mathcal{F})}^{\frac{1}{2}}$
$\lesssim\sum_{Q\in\mathfrak{C}}2^{-\frac{kd}{2}}\bigg(\int_{3Q\setminus\mathcal{F}}\sum_{I\in\mathcal{E}_{k}}\sup_{t\in I}\int_{\mathbf{S}^{2d-1}}\left((|h^{i}_{1}|*\Psi_{2^{-k-n_{1}}})\chi_{{}_{Q}}\right)(x-ty)$
$\hskip 125.19194pt\times\left((|f_{2}|*\Psi_{2^{-k-n_{2}}})\chi_{{}_{5Q}}\right)(x-tz)\;d\sigma(y,z)\;dx\bigg)^{\frac{1}{2}}$
$\lesssim\sum_{Q\in\mathfrak{C}}2^{-\frac{kd}{2}}\bigg(\int_{3Q}\sum_{I\in\mathcal{E}_{k}}\int_{\mathbf{S}^{2d-1}}\left((|h^{i_{1}}_{1}|*\Phi_{2^{-k-n_{1}}})\chi_{{}_{2Q}}\right)(x-t_{0}y)$
$\hskip 125.19194pt\times\left((|h^{i_{2}}_{2}|*\Psi_{2^{-k-n_{2}}})\chi_{{}_{6Q}}\right)(x-t_{0}z)\;d\sigma(y,z)\;dx\bigg)^{\frac{1}{2}}$
$\lesssim 2^{-\frac{kd}{2}}\sum_{Q\in\mathfrak{C}}\bigg(\sum_{I\in\mathcal{E}_{k}}t_{0}^{-d}\big\|(|h^{i_{1}}_{1}|*\Phi_{2^{-k-n_{1}}})\chi_{{}_{2Q}}\big\|_{L^{1}}\;\big\|(|h^{i_{2}}_{2}|*\Psi_{2^{-k-n_{2}}})\chi_{{}_{6Q}}\big\|_{L^{1}}\bigg)^{\frac{1}{2}}$
$\leq 2^{-\frac{kd}{2}}2^{|\mathbf{n}|\frac{\beta}{2}}2^{\frac{kd}{2}}\bigg(\sum_{Q\in\mathfrak{C}}\|(|h^{i_{1}}_{1}|*\Phi_{2^{-n_{1}}})\chi_{{}_{2Q}}\|_{L^{1}}\bigg)^{\frac{1}{2}}\bigg(\sum_{Q\in\mathfrak{C}}\|(|h^{i_{2}}_{2}|*\Psi_{2^{-n_{2}}})\chi_{{}_{6Q}}\|_{L^{1}}\bigg)^{\frac{1}{2}}$
$\lesssim 2^{|\mathbf{n}|\frac{\beta}{2}}2^{(N-d)\frac{i_{1}-k-n_{1}}{2}}\|h^{i_{1}}_{1}\|_{L^{1}}^{\frac{1}{2}}\|h^{i_{2}}_{2}\|_{L^{1}}^{\frac{1}{2}},$
where we have used $L^{1}\times L^{1}\to L^{1}$-boundedness of single scale bilinear spherical average in the third inequality and Hölder’s inequality in the fourth inequality in the expression above.
Since $i_{1}<k$, we have $2^{(N-d)\frac{i_{1}-k-n_{1}}{2}}\leq 2^{\frac{i_{1}-k}{2}}$ for every $N\geq d+1$. We choose $N=d+1$ to obtain
$\|\mathcal{M}^{\mathbf{n}}_{k}(h^{i_{1}}_{1},h^{i_{2}}_{2})\|_{L^{1/2}(\mathbb{R}^{d}\setminus\mathcal{F})}^{\frac{1}{2}}\lesssim 2^{\frac{i_{1}-k}{2}}2^{|\mathbf{n}|\frac{\beta}{2}}\|h^{i_{1}}_{1}\|_{L^{1}}^{\frac{1}{2}}\|h^{i_{2}}_{2}\|_{L^{1}}^{\frac{1}{2}}.$
Interchanging the roles of the $i_{1}$ and $i_{2}$ proves the desired bound for $i_{2}$. This completes the proof of Lemma 6.4. Consequently, the proof of part $(3)$ of Lemma 6.2 is completed. ∎
Acknowledgements :
S.S. Choudhary is supported by the National Center for Theoretical Sciences under project no. 114-2124-M-002-004-. C.-Y. Shen is supported in part by NSTC through grant 111-2115-M-002-010-MY5.